# The evolution of cooperation with Q-learning: the impact of information perception

### v1

@@ Abstract
The inherent huge complexities in human beings show a remarkable diversity in response to complex surroundings, enabling us to tackle problems from different perspectives. In the realm of cooperation studies, however, existing work assumes that individuals get access to the same kind of information to make their decisions, in contrast to the facts that individuals often perceive differently. Here, within the reinforcement learning framework, we investigate the impact of information perception on the evolution of cooperation in a 2-person scenario when playing the prisoner’s dilemma game. We demonstrate that distinctly different evolution processes are observed in three information perception scenarios, revealing that the structure of information significantly affects the emergence of cooperation. Notably, the asymmetric information scenario exhibits a rich dynamical process, including the cooperation emergence, breakdown, and reconstruction, akin to psychological changes in humans. Our findings indicate that the information structure is vital to the emergence of cooperation, shedding new light on establishing mutually stable cooperative relationships and understanding human behavioral complexities in general.
I 1. Introduction
Cooperation is the basis for the survival, development, and reproduction of human beings, and it is ubiquitous in human societies as well as for other species Axelrod and Hamilton (1981a); Maynard Smith and Szathmáry (1995); Jusup et al. (2022). A robust cooperation between individuals effectively improves the work efficiency and benefits the collective. However, due to its complexity and subtlety, non-cooperation frequently occurs, such as global warming, overfishing, deforestation, grazing, war conflicts, and so on, which could be catastrophic. So how cooperation emerges, and under what condition it breaks down? The emergence and maintenance of cooperation have consistently been fundamental challenges for humanity Pennisi (2005).
Evolutionary game theory Nowak and Sigmund (2004); Roca et al. (2009a) has served as the main framework for the study of cooperation, with prototypical models being adopted such as the prisoner’s dilemma (PD) game Doebeli and Hauert (2005). The PD game shows that even when cooperation is beneficial for both parties as a collective, maintaining cooperation remains difficult. This difficulty arises from the fact that individuals seek to maximize their own interests and choose defection, leading to the tragedy of commons Rapoport and Chammah (1965). Thus, there must be some mechanisms that are able to overcome this dilemma to foster cooperation.
@@
Several mechanisms for the emergence of cooperation have been proposed in the past decades Nowak (2006); Perc et al. (2017), including direct Trivers (1971) and indirect reciprocity Nowak and Sigmund (1998), kin and group selection Queller (1964), punishment and reward Sigmund et al. (2001), network Nowak and May (1992); Szabó and Tőke (1998); Wang et al. (2013) and dynamical reciprocity Liang et al. (2022), social diversity Perc and Szolnoki (2008); Santos et al. (2008); Liang et al. (2021), reputation Xia et al. (2023), and behavioral multimodality Ma et al. (2023) etc. Note that these game-theoretic studies typically employ imitation learning Roca et al. (2009b), such as the Moran rule Knight et al. (2018), Fermi updating rule Szabó and Tőke (1998); Szabó et al. (2005), and follow-the-best rule Nowak and Sigmund (1993) et al. The idea behind is that individuals are more likely to imitate strategies of neighbors who are better off, which can be viewed as a simple version of social learning Bandura and Walters (1977).
@@
Reinforcement learning (RL) Sutton and Barto (2018) as an alternative paradigm provides a completely different perspective to study the evolution of cooperation. There, players score different actions within different states, and by repeatedly interacting with the environment they are able to make decisions by balancing the past experience, the present reward, and the expected earnings in the future. Despite its great potential Lee (2008); Silver et al. (2016); Subramanian et al. (2022), RL as a distinct learning paradigm from imitation learning has been largely overlooked. Only recently, researchers start to apply reinforcement learning to evolutionary game theory to help understand the emergence of cooperation Tanabe and Masuda (2012); Ezaki et al. (2016); Horita et al. (2017); Ding et al. (2019); Fan et al. (2022); Zhang et al. (2020); Wang et al. (2022, 2023); He et al. (2022); Ding et al. (2023); Geng et al. (2022); Yang et al. (2024); Shi and Rong (2022); Zhang et al. (2024), trust Zheng et al. (2024), resource allocation Andrecut and Ali (2001); Zhang et al. (2019), and other collective behaviors for humans Tomov et al. (2021); Shi and Rong (2022).
In particular, some new insights have been obtained for the emergence of cooperation within the RL framework. For example, Zhang et al. revealed that explosive cooperation appears in the form of periodic oscillation in snowdrift games with reinforcement learning Zhang et al. (2020). By considering the uncertainties in daily life, Wang et al. uncovered that Lévy noise promotes the evolution of cooperation through reinforcement learning Wang et al. (2022). A year later, they proposed a framework by combining the public goods game with an adaptive reward mechanism within the reinforcement learning framework. The simulations demonstrate that the fraction of cooperation increases significantly when the adaptive reward strategy is included Wang et al. (2023). He et al. extended the PD game to mobile populations, identifying three different types of spontaneous segregation. Adaptive migration enhances network reciprocity and enables the dominance of cooperation in a dense population He et al. (2022). In the two-player scenario, Ding et al. revealed that a strong memory and long-sighted expectation yield the emergence of coordinated optimal policies, where both agents act like “win-stay, lose-shift” to maintain a high level of cooperation Ding et al. (2023). Current game-theoretic studies also indicate the reinforcement learning is able to catalyze cooperation when it is mixed with other updating rules Geng et al. (2022); Han et al. (2022); Sheng et al. (2024). In all of above work, however, they adopt the same state setup, i.e. individuals are assumed to perceive the same kind of information, either the agent’s own action information Zhang et al. (2020); Wang et al. (2022, 2023) or the neighbors’  He et al. (2022), or both Ding et al. (2023); Geng et al. (2022). These information perception is all symmetric in terms of information structure.
Yet, numerous real-world observations indicate that this symmetric perception is often not the case. In fact, factors such as age, experiences, culture, social status, wealth, personality, personal religious beliefs and sources of information directly shape an individual’s unique perspective on issues Dawkins (1989); Molinas (1998). Likewise, indirect factors like economic, social, cultural and political environment Henrich et al. (2004), media influence, history and traditions, globalization, collectively shape societal views on problems, leading to a diverse range of perspectives. As the consequence, players may focus on the different part of the information available to them Kagel et al. (1996); Allison (1978); Feltovich (2000); McAvoy and Hauert (2015), and we have no idea about the consequence of these variations in the information intake. Therefore, it’s natural to ask how different information perceptions affect the evolution of cooperation.
In this work, we focus on the impact of information on the evolution of cooperation within the reinforcement learning paradigm. Specifically, we employ the Q-learning algorithm Watkins (1989, 1992), wherein each individual possesses a Q-table as an action guide and plays the PD game Doebeli and Hauert (2005). By studying three simplest setups of information perception, we uncover different mechanisms of cooperation evolution in the three two-person scenarios. In two of them, the cooperation shows a first-order-like phase transition with the dilemma strength. Interestingly, in the asymmetric information scenario, the time evolution of the cooperation exhibits rich dynamical behaviors, undergoing cooperation emergence – breakdown – reestablishment. As a result, the highest cooperation preference is reached within the shortest time on average in this scenario.
@@
This paper is organized as follows: we introduce our Q-learning model with three different information scheme in Sec. 2II. In Sec. 3III, we present the results. In Sec. 4IV, we provide a mechanistic analysis. In Sec. 5IV, the evolution processes for both symmetric and asymmetric information scenarios are compared. Finally, we conclude our work together with discussions in Sec. 6VI.
@@
II 2. Model
Figure 2: The dependence of cooperation preference on the dilemma strength within the three schemes. (a-c) The time-averaged cooperation preference $\left\langle f_{c}\right\rangle_{t}$ versus the dilemma strength $b$, respectively for Scheme I, II, III. While no clear dependence is observed in Scheme I, the dependence shows a discontinuous transition of cooperation preference in Scheme II and III. The two insets shows typical time series of $f_{c}$ for $b=0.3$ in the corresponding scheme; the red and blue lines represent the results of evolution from two different initial conditions in (b). This means that once the system evolves into mutual cooperation or mutual defection, no change is expected. But persist state switches between the two solutions are always observed in (c). (d-f) The corresponding probability density function (PDF) curve of $f_{c}$, respectively, for Scheme I-III, where trimodal distribution is seen for Scheme I, and bimodal distributions are for the other two schemes. The dashed lines in (e, f) indicate the peak value of $f_{c}=1$ where $b=0.3$, where a higher value is observed in Scheme III than Scheme II. Each data is averaged 500 times after a transient of $3\times 10^{8}$ rounds in (a)-(c). Other parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@
where $R=1.0$, $S=-b$, $T=1+b$, and $P=0$ adopted in our study, corresponding to a strong version of PD Axelrod and Hamilton (1981b). $b>0$ is the dilemma strength, a larger value of which means less likely for cooperation to survive.
@@
Specifically, we consider three different information schemes. (I) Both players are informed of the opponent’s action; (II) Both players consider one’s own action information; (III) One player considers the opponent’s action information, while the other considers one’s own action information in the last round. Obviously, in either Scheme I or II the information used is structurally symmetric for the two players, but this is not the case in Scheme III, where they both concern the action of one player, and is thus asymmetric. The illustration of the three schemes is shown in Fig. 1.
The evolution of the two-player system follows a synchronous updating procedure. At the beginning, each player is randomly assigned an initial strategy C or D as the state, and the elements $Q_{{s_{l}},{a_{m}}}(l,m=1,2)$ in the Q-tables are randomly assigned a value between $[0,1)$, indicating that individuals are initially unfamiliar with the environment. Given the state $s$ at round $t$, (i) with a probability $\epsilon$ each player randomly chooses an action $a\in\mathbb{A}$ to conduct a trial-and-error exploration, otherwise each chooses an action $a$ according to one’s Q-table (i.e. $a$ is selected given $Q_{s,a}>Q_{s,\hat{a}}$). (ii) Then, two players play the PD game and get a payoff $\pi$ according to the matrix Eq. (1). (iii) They get their new state $s^{\prime}$ and update their Q-tables. Specifically, the element $Q_{s,a}(t)$ just referred is updated as follows:
@@
where $\alpha\in(0,1]$ is the learning rate, capturing the contribution at the current step, a larger value of $\alpha$ means that the player is more forgetful as old Q-values tend to be rapidly modified. $\pi(t)$ is the payoff obtained at present round following the payoff matrix Eq. (1). $\gamma\in[0,1)$ is the discount factor, measuring the weight of future rewards, as $\max_{a^{\prime}}Q_{s^{\prime},a^{\prime}}(t)$ is the maximal value expected within the new state. The r.h.s. of the above equation indicates that the Q-values simultaneously contain the contribution of past experiences, reward at present and from the future.
The above process [steps (i)-(iii)] is repeated until the system reaches an equilibrium or the desired duration is completed. The three learning parameters are fixed at typical values of $\epsilon=0.01$, $\alpha=0.1$, $\gamma=0.9$ throughout the study, where the players appreciate both past experiences and expected rewards in the future.
III 3. Results
We report the evolution of cooperation for the three information schemes, where discontinuous transitions and bistability are uncovered, see Fig. 2. As shown in Fig. 2(a), when players focus on the opponent’s action information (Scheme I), cooperation exhibits strong instability even at small values of temptation $b$. With the increase of $b$, the system evolves to a stable state dominated by mutual defection $f_{c}\approx 0$. Correspondingly, the probability density function (PDF) curves of $f_{c}$ within the unstable interval in Fig. 2(d) shows a trimodal distribution. With increasing $b$, the peaks at $0.5$ and $1$ both reduce.
@@
These results suggest that the information structure has a huge impact on the evolution of cooperation, and asymmetric information leads to new complexities in the form of first-order phase transition and true bistability.
@@
IV 4. Mechanism Analysis
Here, we primarily analyze the mechanisms under the asymmetric scenario in Scheme III. The mechanism analyses for Schemes I and II are relatively straightforward and can be found in Appendices B and C, respectively.
@@
    Stage III: how does cooperation re-establishe afterwards?
Figure 4: The action preference combinations of two players within two states. The four quadrants, based on the sign of the value $\Delta Q_{s_{l}}^{i}$ ($i\in\{M,Y\}$), represent the four possible combinations of action preferences. The left and right figures correspond to the system being in state C and state D, respectively. For example, the combination (D, C) in pink indicates that in state D, individual M prefers action D, while individual Y prefers action C.
In addition to the elements $Q^{i}_{s_{l},a_{m}}$ for each player $i$, we are particularly interested in their relative magnitude within a given row, i.e. $\Delta Q_{s_{l}}^{i}=Q_{s_{l},a_{1}}^{i}-Q_{s_{l},a_{2}}^{i}$. This value determines which action is preferred for player $i$ within the given state $s_{l}$. For example, if $\Delta Q_{s_{l}}^{i}\textgreater 0$, this means that for player $i$, the action C is preferred within the state $s_{l}$, otherwise D is supposed to be a better choice. Accordingly, we explicitly show the action preference combinations within two states (see Fig. 4) based on the sign of $\Delta Q_{s_{l}}^{i}$, where $i\in\{M,Y\}$, respectively, labels the one who considers one’s own information of action (“Me”) and the one who considers the opponent’s action information (“You”). For example, action preference combination (D,C) represents that individual M chooses action D and individual Y chooses action C, which in different states are represented by different colors, state C (black), state D (pink).
To be certain, we start with a typical initial condition far from mutual cooperation $Q_{C,C}^{M}<Q_{C,D}^{M}$, $Q_{D,C}^{M}<Q_{D,D}^{M}$, $Q_{C,C}^{Y}>Q_{C,D}^{Y}$, $Q_{D,C}^{Y}<Q_{D,D}^{Y}$ and analyze the dynamical mechanisms.
Figure 5: Cooperation emergence in Stage I. It shows the evolution of action combination preferences, and the temporal evolution of $Q_{s_{l},a_{m}}^{M,Y}$ or $\Delta Q_{s_{l}}^{M,Y}$ values. Here, the action preference combination (C,D)(D,C) indicates that individual M chooses defect in state C and cooperate in state D, causing the system to cycle between (C,D)$\leftrightarrow$(D,C). The same applies to (D,D)(C,D). The sudden declines in (a) are because of occasional cooperation by exploration, where the action of defection brings to a reward $\pi=1+b$. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.

### v2

# Evolution of cooperation with Q-learning: the impact of information perception

@@ Abstract
The inherent complexity of human beings manifests in a remarkable diversity of responses to intricate environments, enabling us to approach problems from varied perspectives. However, in the study of cooperation, existing research within the reinforcement learning framework often assumes that individuals have access to identical information when making decisions, which contrasts with the reality that individuals frequently perceive information differently. In this study, we employ the Q-learning algorithm to explore the impact of information perception on the evolution of cooperation in a two-person Prisoner’s Dilemma game. We demonstrate that the evolutionary processes differ significantly across three distinct information perception scenarios, highlighting the critical role of information structure in the emergence of cooperation. Notably, the asymmetric information scenario reveals a complex dynamical process, including the emergence, breakdown, and reconstruction of cooperation, mirroring psychological shifts observed in human behavior. Our findings underscore the importance of information structure in fostering cooperation, offering new insights into the establishment of stable cooperative relationships among humans.
preprint: AIP/123-QED
In the real world, our perceptions of information are shaped by a variety of factors, leading to diverse responses to environmental stimuli and underscoring the importance of perceptual differences in decision-making processes. To explore how these differences influence the evolution of cooperation, we develop a simplified two-player Prisoner’s Dilemma model using the Q-learning algorithm. By analyzing three distinct information perception scenarios, we observe significantly different evolutionary processes, with the asymmetric information scenario exhibiting particularly complex dynamics in the emergence and stability of cooperation. These findings emphasize the critical role of information structure in shaping cooperative behaviors and provide new insights into the complexities of human decision-making.
I Introduction
Cooperation is fundamental to the survival, development, and reproduction of humans and other species, playing a crucial role in improving collective efficiency and benefits Axelrod and Hamilton (1981a); Maynard Smith and Szathmáry (1995); Jusup et al. (2022). However, its complexity and subtlety often lead to non-cooperation, manifesting in issues like global warming, overfishing, and conflicts, which can have catastrophic consequences. Understanding how cooperation emerges and under what conditions it breaks down remains a central challenge Pennisi (2005). Evolutionary game theory Szabó and Fáth (2007); Perc and Szolnoki (2010), particularly through models like the prisoner’s dilemma (PD) game Rapoport and Chammah (1965), has been instrumental in studying cooperation. The PD game illustrates the difficulty of maintaining cooperation despite its collective benefits, as individuals tend to prioritize self-interest and defect. Identifying mechanisms that overcome this dilemma to promote cooperation is therefore essential.
@@ I Introduction
Several mechanisms for the emergence of cooperation have been proposed in the past decades Nowak (2006); Perc et al. (2017), including direct Trivers (1971) and indirect reciprocity Nowak and Sigmund (1998), kin and group selection Queller (1964), punishment and reward Sigmund, Hauert, and Nowak (2001), network Nowak and May (1992); Szabó and Tőke (1998); Wang, Szolnoki, and Perc (2013) and dynamical reciprocity Liang et al. (2022), social diversity Perc and Szolnoki (2008); Santos, Santos, and Pacheco (2008); Liang et al. (2021), reputation Xia et al. (2023), and behavioral multimodality Ma et al. (2023) etc. Note that these game-theoretic studies typically employ imitation learning Roca, Cuesta, and Sánchez (2009), such as the Moran rule Knight et al. (2018), Fermi-function-based update rule Szabó and Tőke (1998); Szabó, Vukov, and Szolnoki (2005), and follow-the-best rule Nowak and Sigmund (1993) et al. The idea behind is that individuals are more likely to imitate strategies of neighbors who are better off, which can be regarded as a simplified form of social learning Bandura and Walters (1977).
@@
Reinforcement learning (RL) Sutton and Barto (2018) as an alternative paradigm provides a fundamentally different perspective to study the evolution of cooperation Wang, Fu, and Chen (2024). In RL, players score different actions within different states, and by repeatedly interacting with the environment they are able to make decisions by balancing the past experience, the present reward, and the expected earnings in the future. Despite its great potential Lee (2008); Silver et al. (2016); Subramanian, Chitlangia, and Baths (2022), RL as a distinct learning paradigm from imitation learning has been largely overlooked. Recently, researchers have started to apply reinforcement learning to evolutionary game theory to help understand the evolution of social behaviors, such as cooperation Tanabe and Masuda (2012); Fan et al. (2022); Shi and Rong (2022); Zhang et al. (2020); Wang et al. (2022, 2023a); He et al. (2022); Ding et al. (2023); Geng et al. (2022); Zhang et al. (2024); Zheng et al. (2024a); Mintz and Fu (2025), trust Zheng et al. (2024b), fairness Zheng et al. (2024c), collective motion López-Incera et al. (2020); Wang et al. (2023b), and resource allocation Andrecut and Ali (2001); Zhang et al. (2019).
This growing body of work highlights the versatility of RL in understanding complex social dynamics and its potential to uncover new insights into the mechanisms driving cooperative and collective behaviors. For instance, Zhang et al. demonstrated that explosive cooperation manifests as periodic oscillations in snowdrift games using RL Zhang et al. (2020). Wang et al. found that Lévy noise enhances cooperation through RL, accounting for real-world uncertainties Wang et al. (2022). Later, they integrated an adaptive reward mechanism into the public goods game, showing a significant increase in cooperation levels Wang et al. (2023a). He et al. extended the PD game to mobile populations, revealing that adaptive migration strengthens network reciprocity and promotes cooperation in dense populations He et al. (2022). In two-player scenarios, Ding et al. showed that coordinated optimal policies emerge from strong memory and long-term expectations, with agents adopting a "win-stay, lose-shift" strategy to sustain high cooperation Ding et al. (2023). Additionally, studies suggest that RL can catalyze cooperation when combined with other updating rules Geng et al. (2022); Han, Zhao, and Xia (2022); Sheng et al. (2024). However, these works assume symmetric information perception, where individuals access the same type of information, such as their own actions Zhang et al. (2020); Wang et al. (2022, 2023a), neighbors’ actions He et al. (2022), or both Ding et al. (2023); Geng et al. (2022).
Yet, numerous real-world observations indicate that information perception is often asymmetric, shaped by factors like age, experience, culture, social status, and personal beliefs Dawkins (1989); Molinas (1998), as well as indirect influences such as economic, social, and political environments Henrich et al. (2004). This diversity leads individuals to focus on different aspects of available information Kagel, Kim, and Moser (1996); Allison (1978); Feltovich (2000); McAvoy and Hauert (2015), raising the question of how such variations in information perception impact cooperation. While some studies highlight the role of information richness in cooperation Jia et al. (2021); Yang et al. (2024); Song et al. (2022), they often rely on network structures and neighbor payoff information, leaving the more fundamental pairwise interactions and action-based information unexplored.
In this work, we adopt a fresh perspective by distinguishing between information sources within the RL framework, focusing on action information rather than payoffs. Using the Q-learning algorithm Watkins (1989, 1992), we systematically investigate cooperation evolution under symmetric and asymmetric information settings in two-player PD games Doebeli and Hauert (2005). We identify distinct mechanisms across three information perception scenarios, revealing rich dynamical behaviors in the asymmetric case, including cooperation emergence, breakdown, and reestablishment. Notably, the asymmetric scenario achieves the highest cooperation preference in the shortest time, underscoring the critical role of information structure in shaping cooperative dynamics.
@@
This paper is organized as follows: we introduce our Q-learning model with three different information schemes in Sec. II. In Sec. III, we present the results. In Sec. IV, we provide a mechanistic analysis. In Sec. V, the evolution processes for both symmetric and asymmetric information scenarios are compared. Finally, we conclude our work together with discussions in Sec. VI.
II Model
@@
where $R=1.0$, $S=-b$, $T=1+b$, and $P=0$ are adopted in our study, corresponding to a strong version of PD Axelrod and Hamilton (1981b). $b>0$ is the dilemma strength, a larger value of which means less likely for cooperation to survive. A more general understanding of dilemma strength in symmetric 2×2 games can be found in refs. Tanimoto and Sagara (2007); Ito and Tanimoto (2018).
Figure 2: The dependence of cooperation preference on the dilemma strength within the three schemes. (a-c) The time-averaged cooperation preference $\left\langle f_{c}\right\rangle_{t}$ versus the dilemma strength $b$, respectively for Scheme I, II, III. While no clear dependence is observed in Scheme I, the dependence shows a discontinuous transition of cooperation preference in Scheme II and III. The two insets show typical time series of $f_{c}$ for $b=0.3$ in the corresponding scheme; the red and blue lines represent the results of evolution from two different initial conditions in (b). This means that once the system evolves into mutual cooperation or mutual defection, no change is expected. But persist state switches between the two solutions are always observed in (c). (d-f) The corresponding probability density function (PDF) curve of $f_{c}$, respectively, for Scheme I-III, where trimodal distribution is seen for Scheme I, and bimodal distributions are for the other two schemes. The dashed lines in (e, f) indicate the peak value of $f_{c}=1$ where $b=0.3$, where a higher value is observed in Scheme III than Scheme II. Each data is averaged 500 times after a transient of $3\times 10^{8}$ rounds in (a)-(c). Other parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@
Specifically, we consider three different information schemes. (I) Both players are informed of the opponent’s action; (II) Both players consider one’s own action information; (III) One player considers the opponent’s action information, while the other considers one’s own action information in the last round. Obviously, in either Scheme I or II, the information used is structurally symmetric for the two players, but this is not the case in Scheme III, where they both concern the action of one player, and is thus asymmetric. The illustration of the three schemes is shown in Fig. 1.
The evolution of the two-player system follows a synchronous updating procedure. At the beginning, each player is randomly assigned an initial strategy C or D as the state, and the elements $Q_{{s_{l}},{a_{m}}}(l,m=1,2)$ in the Q-tables are randomly assigned a value between $(0,1)$, indicating that individuals are initially unfamiliar with the environment. At round $t$, given the state $s$: (i) With a probability $\epsilon$, each player randomly chooses an action $a\in\mathbb{A}$ to conduct trial-and-error exploration; otherwise, each chooses an action $a$ according to one’s Q-table (i.e., $a$ is selected if $Q_{s,a}>Q_{s,\hat{a}}$). (ii) Then, two players play the PD game and get a payoff $\pi$ according to the matrix Eq. (1). (iii) They get their new state $s^{\prime}$ and update their Q-tables. Specifically, the element $Q_{s,a}(t)$ just referred is updated as follows:
@@
where $\alpha\in(0,1]$ is the learning rate, which captures the contribution of the current step. A larger $\alpha$ means that the player is more forgetful, as old Q-values tend to be more rapidly modified. $\pi(t)$ is the payoff obtained at present round following the payoff matrix Eq. (1). $\gamma\in[0,1)$ is the discount factor, measuring the weight of future rewards, as $\max_{a^{\prime}}Q_{s^{\prime},a^{\prime}}(t)$ is the maximal value expected within the new state. The r.h.s. of the above equation indicates that the Q-values simultaneously contain the contribution of past experiences, reward at present and from the future.
The above process [steps (i)-(iii)] is repeated until the system reaches an equilibrium or the desired duration is completed. The three learning parameters are fixed at typical values of $\epsilon=0.01$, $\alpha=0.1$, $\gamma=0.9$ throughout the study, where players appreciate both past experiences and expected rewards in the future.
III Results
We report the evolution of cooperation for the three information schemes, where discontinuous transitions and bistability are uncovered, see Fig. 2. As shown in Fig. 2(a), when players focus on the opponent’s action information (Scheme I), cooperation exhibits strong instability even at small values of temptation $b$. With the increase of $b$, the system evolves to a stable state dominated by mutual defection $f_{c}\approx 0$. Correspondingly, the probability density function (PDF) curves of $f_{c}$ within the unstable interval in Fig. 2(d) show a trimodal distribution. With increasing $b$, the peaks at $0.5$ and $1$ both reduce.
@@ III Results
These results suggest that the information structure has a huge impact on the evolution of cooperation, and asymmetric information leads to new complexities in the form of first-order-like phase transition and true bistability.
@@
IV Mechanism Analysis
Figure 4: The action preference combinations of two players within two states. The four quadrants, based on the sign of the value $\Delta Q_{s_{l}}^{i}$ ($i\in\{M,Y\}$), represent the four possible combinations of action preferences in different states, denoted with subscripts. The left and right figures correspond to the system being in state C and state D, respectively. For example, the combination $(D,C)_{D}$ indicates that in state D, individual M prefers action D, while individual Y prefers action C.
Here, we primarily analyze the mechanisms under the asymmetric scenario in Scheme III. The mechanism analyses for Schemes I and II are relatively straightforward and are provided in Appendices B and  C, respectively.
Figure 5: Cooperation emergence in Stage i. It shows the evolution of action combination preferences, and the temporal evolution of $Q_{s_{l},a_{m}}^{M,Y}$ or $\Delta Q_{s_{l}}^{M,Y}$ values. Here, the action preference combination $(C,D)_{D}$$(D,C)_{C}$ indicates that individual M chooses defect in state C and cooperate in state D, causing the system to cycle between $(C,D)_{D}$$\leftrightarrow$$(D,C)_{C}$. The same applies to $(D,D)_{C}$$(C,D)_{D}$. The sudden declines in (a) are because of occasional cooperation by exploration, where the action of defection brings to a reward $\pi=1+b$. Parameters: $\epsilon$ = 0.01, $\alpha$ = 0.1, $\gamma$ = 0.9.
@@ IV Mechanism Analysis
    Stage iii: how does cooperation reestablish afterwards?
In addition to the elements $Q^{i}_{s_{l},a_{m}}$ for each player $i$, we are particularly interested in their relative magnitude within a given row, i.e., $\Delta Q_{s_{l}}^{i}=Q_{s_{l},a_{1}}^{i}-Q_{s_{l},a_{2}}^{i}$. This value determines which action is preferred for player $i$ within the given state $s_{l}$. For example, if $\Delta Q_{s_{l}}^{i}\textgreater 0$, this means that for player $i$, the action C is preferred within the state $s_{l}$, otherwise D is supposed to be a better choice. Accordingly, we explicitly show the action preference combinations within two states [see Fig. 4] based on the sign of $\Delta Q_{s_{l}}^{i}$, where $i\in\{M,Y\}$ labels the individual who considers their own action information (“Me”) and the individual who considers the opponent’s action information (“You”). For example, the action preference combination $(D,C)$ represents individual M choosing action D and individual Y choosing action C, which are denoted by different subscripts in different states: state C is represented by $(D,C)_{C}$, and state D by $(D,C)_{D}$.
To be certain, we start with a typical initial condition that is far from mutual cooperation $Q_{C,C}^{M}<Q_{C,D}^{M}$, $Q_{D,C}^{M}<Q_{D,D}^{M}$, $Q_{C,C}^{Y}>Q_{C,D}^{Y}$, $Q_{D,C}^{Y}<Q_{D,D}^{Y}$ and analyze the dynamical mechanisms. For other cases of total betrayal, refer to the evolution process in Stage iii.